# Light and strange vector resonances from lattice QCD at physical quark masses

### v1

@@ Acknowledgements.
The authors thank the members of the RBC and UKQCD Collaborations for the helpful discussions and suggestions. N.P.L., F.E. and A.P. kindly thank Mike Peardon for the invaluable discussions. N.P.L. additionally thanks André Baião Raposo for the discussions. This work used the DiRAC Extreme Scaling service (Tursa / Tesseract) at the University of Edinburgh, managed by the Edinburgh Parallel Computing Centre on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Edinburgh was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. P.B. has been supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under the Contract No. DE-SC-0012704 (BNL). P.B. has also received support from the Royal Society Wolfson Research Merit award WM/60035. M.M. gratefully acknowledges support from STFC in the form of a fully funded PhD studentship. A.P. & F.E. received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No 757646. N.P.L. & A.P. received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No 813942. F.E. has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101106913. M.T.H. and F.J. are supported by UKRI Future Leader Fellowship MR/T019956/1. A.P., M.T.H., V.G. & F.E. were supported in part by UK STFC grant ST/P000630/1, and A.P., M.T.H. & V.G. additionally by UK STFC grants ST/T000600/1 & ST/X000494/1.

### v2

@@ Abstract
Acknowledgments—The authors thank the members of the RBC and UKQCD Collaborations for the helpful discussions and suggestions. N. P. L., F. E. and A. P. kindly thank Mike Peardon for the invaluable discussions. N. P. L. additionally thanks André Baião Raposo for the discussions. This work used the DiRAC Extreme Scaling service (Tursa / Tesseract) at the University of Edinburgh, managed by the Edinburgh Parallel Computing Centre on behalf of the STFC DiRAC HPC Facility. The DiRAC service at Edinburgh was funded by BEIS, UKRI and STFC capital funding, and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure. P. B. has been supported in part by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under the Contract No. DE-SC-0012704 (BNL). P. B. has also received support from the Royal Society Wolfson Research Merit award WM/60035. M. M. gratefully acknowledges support from STFC in the form of a fully funded Ph.D. studentship. A. P. and F. E. received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 757646. N. P. L. and A. P. received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 813942. N. P. L. also acknowledges support from the United Kingdom Science and Technology Facilities Council (STFC) [Grants No. ST/T000694/1, ST/X000664/1]. F. E. has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101106913. M. T. H. and F. J. are supported by UKRI Future Leader Fellowship No. MR/T019956/1. A. P., M. T. H., V. G. and F. E. were supported in part by United Kingdom STFC Grant No. ST/P000630/1, and No. A. P., M. T. H. and V. G. additionally by UK STFC grants No. ST/T000600/1 and No. ST/X000494/1. A. P. is additionally partially supported by a long-term Invitational Fellowship from the Japan Society for the Promotion of Science.